OpenStudy (anonymous):

how do you find the zeros of the function f(x)=x(x-1)(x-5)^3

OpenStudy (phi):

you set that mess equal to zero: x(x-1)(x-5)^3 =0 now use this idea: if any of the 3 terms (x or (x-1) or (x-5)^3 is zero) then the whole thing is zero. Think about that. That means you can break this into 3 separate problems: x= 0 x-1 = 0 (x-5)^3 = 0

OpenStudy (anonymous):

thank you, that was VERY helpful!

OpenStudy (phi):

I should add that (x-5)^3 = 0 can be written as (x-5)(x-5)(x-5) =0 which turns into 3 (identical) problems: x-5 =0 x-5= 0 x-5 =0 so all to0gther there are 5 separate equations x= 0 x-1 = 0 x-5 =0 x-5= 0 x-5 =0 and 5 solutions: 0, 1 , 5, 5, 5 the 5's are called "repeated roots"